# Assessing Small Language Models for Code Generation: An Empirical Study with Benchmarks

### v3

@@ 7 Conclusion
Our study shows that using the larger SLMs is not the only way to gain strong performance in code generation tasks. Some smaller and mid-sized SLMs performed surprisingly well. This observation indicates that careful model design and training can improve the models performance, even without scaling up its size to higher parameters.
@@
Our results indicate that there are differences in how models perform across various programming languages. However, those differences are not statistically significant. This suggests that SLMs are generally capable of handling code generation tasks across different programming languages with reasonable consistency.
@@ Acknowledgments
This work has been supported by FAST, the Finnish Software Engineering Doctoral Research Network, funded by the Ministry of Education and Culture, Finland.

### v4

@@ 7 Conclusion
Our results show that strong performance in code generation is not limited to the largest SLMs. Several smaller and mid-sized models (Groups 1 and Group 2, with parameter sizes 3B) performed competitively, in some cases matching or exceeding the performance of larger SLMs (Group 3, with parameter sizes 3B - 10B) within our evaluation range. This suggests that thoughtful model design and training can achieve substantial performance gain without necessarily scaling up the models size to higher parameters.
@@
Our results show that although models vary in performance across programming languages, these differences are not statistically significant. This indicates reasonably consistent behaviour within the languages included in our benchmarks. However, this should not be interpreted as broad multilingual generalisation, because our evaluation covers only a limited set of languages, and the findings apply only to those languages we tested.
@@ Acknowledgments
This work has been partially supported by FAST, the Finnish Software Engineering Doctoral Research Network, funded by the Ministry of Education and Culture, Finland.